### v2

# Fully Parallelized BP Decoding for Quantum LDPC Codes Can Outperform BP-OSD

Abstract.
In this work, we propose a lightweight decoder based solely on belief-propagation (BP), augmented with a speculative post-processing strategy inspired by classical Chase decoding. Our method identifies unreliable bits via BP oscillation statistics, generates a set of modified test patterns, and decodes them in parallel using low-iteration BP. We demonstrate that our approach can achieve logical error rates comparable to or even better than BP-OSD, but has lower latency over its parallelization for a variety of bivariate bicycle codes, which significantly reduces decoding complexity.
1. Introduction
Quantum low-density parity-check (qLDPC) codes have attracted significant attention in recent years due to their potential to encode more logical qubits than surface codes while maintaining a high threshold (Tillich and Zémor, 2014; Panteleev and Kalachev, 2021; Bravyi et al., 2024). However, decoding qLDPC codes remains a major challenge. To ensure reliable operation, especially in fault-tolerant quantum memory and computation, errors must be corrected both quickly and accurately, placing stringent demands on the decoding algorithm in terms of performance and computational efficiency.
Belief propagation (BP)-based decoders, widely used in classical LDPC codes, are appealing due to their low complexity, parallelizability, and near-optimal performance. However, their effectiveness diminishes significantly when applied to qLDPC codes. This degradation is primarily due to inherent properties of qLDPC codes, such as degeneracy, the presence of many low-weight stabilizers, and the prevalence of trapping sets (Raveendran and Vasić, 2021), which hinder convergence and reliability of BP decoding.
Many works have sought to address the limitations of BP-based decoders on qLDPC codes. In (Poulin and Chung, 2008), Poulin and Chung proposed several techniques to enhance convergence, including random freezing of variable nodes, perturbing priors, and colliding unsatisfied check nodes. In (Chytas et al., 2024), instead of performing static trapping set analysis prior to decoding, the authors identify bits affected by trapping sets dynamically during the decoding process. Once identified, the priors of these bits are modified, similar to the approach in (Poulin and Chung, 2008), to help the decoder escape local minima and converge. Raveendran et al. (Raveendran and Vasić, 2021) analyzed different types of trapping sets and proposed using a layered BP decoder to mitigate the effect of symmetric trapping sets. While this approach can reduce complexity, it often comes at the cost of increased decoding latency. More recently, Yin et al. (Yin et al., 2024) leveraged the degeneracy property of quantum LDPC codes to enhance BP decoding. By analyzing bit-wise marginal probabilities from BP, they selectively split rows in the parity-check matrix and modify the corresponding Tanner graph. This symmetry-breaking technique helps the decoder avoid convergence stalls caused by structural degeneracies in the code.
Apart from efforts aimed at improving the BP decoder itself, several works focus on post-processing techniques to enhance decoding performance. A widely used approach is belief propagation combined with ordered statistics decoding (BP-OSD) (Panteleev and Kalachev, 2021; Roffe et al., 2020). While BP-OSD significantly enhances error correction performance, its reliance on a Gaussian elimination step during the OSD phase introduces substantial computational overhead. Specifically, this step incurs a complexity of $O(N^{3})$, where $N$ denotes either the code length (in the code-capacity error model) or the number of error mechanisms (in the circuit-level noise model). In contrast, each iteration of BP has only $O(N)$ complexity. This difference makes BP-OSD computationally expensive and less suitable for large-scale or real-time decoding applications. To address this limitation, recent works such as (Wolanski and Barber, 2024; Hillmann et al., 2024) propose partitioning the Tanner graph into several clusters. This localized decoding approach reduces the size of the matrices involved in Gaussian elimination, thereby lowering the overall computational burden without substantially compromising decoding performance.
In this work, rather than aiming to improve BP itself, we focus on achieving decoding performance comparable to BP-OSD while avoiding computationally expensive post-processing steps such as Gaussian elimination. To this end, we propose a lightweight post-processing technique based entirely on BP and demonstrate that, with appropriate design and a speculative decoding strategy, it can closely match the logical error rates achieved by BP-OSD. Through simulations on various bivariate bicycle codes, we show that the proposed decoder performs similarly in terms of syndrome analysis to BP-OSD with a combination-sweep of order 10. Additionally, the proposed BP decoder requires lower latency given it is fully parallelizable, which means it can outperform BP-OSD in execution efficiency, making it a promising candidate for scalable and efficient decoding of quantum LDPC codes in practical fault-tolerant quantum computing systems.
Throughout the paper, we adopt the widely used min-sum variant of BP because of its simplicity and computational efficiency. Nonetheless, our approach could potentially benefit from incorporating more advanced BP-based techniques.
2. Background
2.1. Quantum LDPC Codes
Stabilizer codes are among the most commonly used codes in quantum error correction. One can measure each stabilizer to infer both the type and location of errors in a multi-qubit system. To construct such a code, all stabilizers must commute with each other. Thus, they have a common eigenspace and form a stabilizer group $\mathcal{S}$. The code space $\mathcal{C}$ defined by such group is
(1) $\mathcal{C}=\{\ket{\psi}|\>s\ket{\psi}=\ket{\psi},\>\forall s\in\mathcal{S}\}.$
An $\llbracket n,k,d\rrbracket$ stabilizer code can be defined by $n-k$ independent stabilizers, allowing us to encode $k$ qubits of logical information into an $n$-qubit block tolerating up to ${\lfloor(d-1)/2\rfloor}$ errors. CSS codes are an important class of stabilizer with two sets of stabilizers, $X$-type and $Z$-type, represented by parity-check matrices $H_{X}$ and $H_{Z}$, respectively. Each row in a parity-check matrix corresponds to a stabilizer generator, and each column corresponds to a physical qubit. A “1” entry indicates an $X$ or $Z$ operator (depending on whether it is in $H_{X}$ or $H_{Z}$), while a “0” indicates the identity. Consequently, an $X$-type stabilizer acts as $X$ or the identity on each qubit, and a $Z$-type stabilizer acts as $Z$ or the identity on each qubit. Errors can therefore be corrected by handling $Z$ errors and $X$ errors separately. Since all stabilizers must commute with each other, it follows directly that for a CSS code $H_{X}H_{Z}^{T}=0$.
2.2. Decoding Problem
Assuming the noise is uniform on each bit, the optimal syndrome decoding problem for classical codes can be formalized as follows: Given a code with parity-check matrix $H\in\mathbb{F}_{2}^{M\times N}$ and syndrome $\bm{s}\in\mathbb{F}_{2}^{M}$, we want to find an error $\hat{\bm{e}}\in\mathbb{F}_{2}^{N}$ that satisfies the syndrome
(2) $\hat{\bm{e}}=\arg\min_{\hat{\bm{e}}H^{T}=\bm{s}}(\sum_{i}^{N}\hat{\bm{e}_{i}}).$
Quantum stabilizer codes face a problem due to the phenomenon of degeneracy, where multiple errors have the same effect on the code space. As a result, the goal of decoding is not to identify the most likely error itself, but rather the most likely equivalence class of errors modulo stabilizers, since errors that differ by a stabilizer operation act identically on the code space. Given a syndrome, $\bm{s}$, the optimal decoding problem considering degeneracy becomes
(3) $[\hat{E}]=\arg\max_{[E]:\text{synd}(E)=\bm{s}}\Pr([E]),$
where $[E]$ denotes the equivalence class of errors under the stabilizer group and $\Pr([E])$ is the total probability of all errors in that class.
While the classical decoding problem in Eq. (2) is NP-hard, its quantum counterpart in Eq. (3) is even more complex, being #P-complete. In practice, optimal decoding is computationally intractable, so efficient decoders typically aim to approximate the solution to Eq. (2) with good performance under realistic noise models.
As CSS codes can be decoded separately on the $X$- and $Z$-error bases, each decoding problem can be treated as a classical decoding task and addressed using classical decoding algorithms. For example, BP, the most commonly used decoder for classical LDPC codes, is also widely applied in qLDPC codes. Given an $M\times N$ parity-check matrix $H$, let $v_{1},\dots,v_{N}$ denote the variable nodes (corresponding to the columns of $H$) and $c_{1},\dots,c_{M}$ the check nodes (corresponding to the rows). The normalized min-sum algorithm, a widely used variant of BP can be described as follows:
  1. (1)
    Initialization: Given the prior error information, $\bm{p}$, of each variable node, $v_{i}$, the channel LLR is initialized to
    (4) $l^{ch}_{v_{i}}=\log\frac{1-p_{v_{i}}}{p_{v_{i}}}.$
  2. (2)
    Variable-to-Check (V2C) Message Update: Each variable node, $v_{i}$, sends a message to its neighboring check node, $c_{j}$, based on the channel LLR and the incoming messages from all other neighboring check nodes, denoted as
    (5) $l_{v_{i}\rightarrow c_{j}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})\setminus\{c_{j}\}}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$
    where $l_{c_{j^{\prime}}\rightarrow v_{i}}$ is set to 0 for the first iteration, $\setminus$” is set minus, and $N(v_{i})$ is the set of all the check nodes connected with $v_{i}$.
  3. (3)
    Check nodes to variable nodes (C2V) update: Each check node, $c_{j}$, updates its message to a neighboring variable node, $v_{i}$, using the min-sum rule denoted as
    (6) $l_{c_{j}\rightarrow v_{i}}=(-1)^{s_{j}}\cdot\alpha\min_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}|l_{v_{i^{\prime}}\rightarrow c_{j}}|\cdot\prod_{v_{i^{\prime}}\in N(c_{j})\setminus\{v_{i}\}}\text{sign}(l_{v_{i^{\prime}}\rightarrow c_{j}}),$
    where $\alpha$ is the damping factor used to attenuate the c2v message.
  4. (4)
    Hard decision: After a fixed number of iterations or upon convergence, the final marginal LLR for each variable node is computed as
    (7) $l^{out}_{v_{i}}=l^{ch}_{v_{i}}+\sum_{c_{j^{\prime}}\in N(v_{i})}l_{c_{j^{\prime}}\rightarrow v_{i}}\>\>,$
    where the $\bm{l}^{out}$ is the marginalized LLR for each variable node. The estimated error is then obtained via hard decision as
    (8) $\hat{e_{i}}=\begin{cases}0&\text{if}\quad l_{v_{i}}^{out}>0\\ 1&\text{otherwise}\end{cases},$
In each iteration, steps 2, 3 and 4 are performed. The BP decoding algorithm proceeds until $\bm{e}H^{T}=\bm{s}$ is satisfied or the maximum number of iterations is reached.
In the classical setting, BP performs close to optimal. However, for qLDPCs, the performance of BP degrades significantly: The decoder may fail to converge due to the high degeneracy of quantum codes and the absence of soft information for each bit, which is typically available in classical scenarios.
3. BP Behavior Analysis: A Case Study
In this section, we analyze the behavior of BP-based decoders on qLDPC codes using the $\llbracket 144,12,12\rrbracket$ “gross” code from (Bravyi et al., 2024) as a representative case study. While specific to this code, the analysis provides general insights into the behavior of BP decoding on qLDPC codes and offers guidance for improving decoder performance. The BP decoder used below is a min-sum decoder with adaptive damping factor $\alpha=1-2^{i}$, where $i$ is the current number of iterations.
3.1. Number of Iterations
Figure 1 shows the non-convergence rate of the min-sum decoder under the circuit-level noise model, where $p$ denotes the physical error rate. The curves are obtained by simulating 10,000 samples for each of $p=0.001$ and $p=0.002$, both representative values below the decoding threshold. For each sample, we record the number of iterations required for convergence and compute the cumulative distribution. The non-convergence rate at iteration $i$ is defined as the fraction of samples that have not converged within $i$ iterations—that is, $1$ minus the cumulative convergence rate.
As Fig. 1 indicates, in most cases, BP converges within a small number of iterations. For instance, at $p=0.001$, the average number of iterations is merely 8.9. Even at higher error rates, such as $p=0.002$, the average number of iterations remains low, although the tail becomes longer. Notably, cases that do not converge within the early iterations rarely benefit from increasing the iteration count further. This observation motivates an alternative strategy: Rather than extending the number of BP iterations, we can vary the inputs to the BP decoder while keeping the maximum number of iterations small. If these varied inputs have better independent chances of successful decoding, then running multiple instances in parallel allows us to exponentially suppress the logical error rate without incurring significant decoding latency.
Figure 1. Ratio of unsuccessful BP decoding ($1-$convergence rate) on the $\llbracket 144,12,12\rrbracket$ code under the circuit-level noise model. The maximum number of decoder iterations is set to 1,000 and number of samples is 10,000.
3.2. Oscillation
Figure 2. Precision and recall probabilities of candidate bit selection on the $\llbracket 144,12,12\rrbracket$ code. We evaluate the correlation between candidate bits and actual error locations by identifying the top 50 most frequently flipped bits among approximately 8,000 error mechanisms. The decoder is run with a maximum of 50 iterations, and statistics are collected over 1,000 decoding failures. This analysis reveals how well bit-level oscillation can serve as a heuristic for error localization.
As suggested in previous works (Raveendran and Vasić, 2021; Chytas et al., 2024), trapping sets and code degeneracy often result in ambiguous BP decoding, leading to convergence failures. A common symptom of such a failure is bit-level oscillation, where certain output bits repeatedly flip between 0 and 1 across iterations. To better understand the relationship between oscillating bits and decoding errors, we analyze the dynamics of bit oscillations during the BP process. During the decoding process, we track bit-level oscillations by comparing the output of each iteration with that of the previous one and counting how often each bit flips, which is similar to Ref (Chytas et al., 2024). We then identify a set of oscillating bits, denoted by $\Phi$, based on their flip frequency. Specifically, $\Phi$ is defined as the top $|\Phi|$ of the most frequently flipped bits. We denote $\operatorname{supp}(\bm{e})$ as the set of erroneous bits and define the hit precision and recall as
(9) Precision $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\Phi|},$
(10) Recall $=\frac{|\operatorname{supp}(\bm{e})\cap\Phi|}{|\operatorname{supp}(\bm{e})|}.$
Fig. 2 shows the precision and recall rate when the min-sum decoder fails to decode a syndrome. We can see that even when the BP decoder fails to fully correct an error, the pattern of bit oscillations essentially reveals a meaningful subset of the actual error locations. In particular, at lower physical error rates, the set of oscillating bits nearly covers the entire true error positions. To confirm, we observe that the hit precision, i.e., the fraction of oscillating bits that are indeed erroneous, is substantially higher than the physical error rate, indicating that this method significantly outperforms random guessing. Therefore, this information can serve as a valuable indicator for identifying unreliable bits during post-processing. As the physical error rate increases, the recall decreases, primarily because the total number of errors grows while the candidate set size remains fixed.
4. Proposed Speculative Decoding Method
To enhance the performance of BP decoding, we adopt a Chase-like post-processing technique (Chase, 1972). This approach generates a set of test vectors by flipping candidate bits and attempts decoding on each of them, thereby increasing the likelihood of successful error correction. In the quantum decoding setting, we only have access to the syndrome rather than the prior probability of each bit being in error. Therefore, unlike the original Chase algorithm, which relies on prior channel information to select candidate bits, our method identifies candidate bits based on BP flipping statistics. As shown in Figure 3, once these candidate bits $\Phi$ are identified, we generate diverse decoding attempts by flipping the input syndrome accordingly across multiple BP instances. This strategy increases the variety in the decoder’s inputs and distinguishes our approach from that in (Koutsioumpas et al., 2025), which does not modify the input syndrome. If the decoder successfully converges on this modified input, we then flip these bits back in the output. This restoration ensures that the final output error matches the original syndrome preserving the validity of the decoding result. Since these candidate bits are likely to correspond to actual error locations, flipping them can effectively reduce the number of errors in the input and equivalently lowers the physical error rate. This reduction not only increases the likelihood of successful decoding but also reduces the number of iterations needed for BP to converge. The detailed pseudo code can be found in Algorithm 1.
Figure 3. A simplified flowchart of the proposed decoder. The full procedure is described in Algorithm 1.
Input : Syndrome $\bm{s}$, max flip weight $w_{\max}$, number of uncertain bits $|\Phi|$
Result: Estimated error $\hat{\bm{e}}$
Function Main($\bm{s},w_{\max},|\Phi|$):
       /* Initial BP attempt with oscillation tracking */
       $s_{\text{ucc}},\hat{\bm{e}},\Phi\leftarrow\textnormal{{BP\_with\_oscillation}}(\bm{s},|\Phi|)$
      if $s_{\text{ucc}}$ then
             return $\hat{\bm{e}}$ /* Syndrome decoded successfully */
      else
             /* Speculative decoding using test patterns based on $\Phi$/
             parallel for $\bm{t}\in\textnormal{{combinations}}(\Phi,w_{\max})$
                   $\bm{s}^{\prime}=\bm{s}\oplus\bm{t}H^{T}$ /* Flip selected bits in syndrome domain */
                   $s_{\text{ucc}},\hat{\bm{e}}\leftarrow\textnormal{{BP}}(\bm{s}^{\prime})$
                   if $s_{\text{ucc}}$ then
                         return $\hat{\bm{e}}\oplus\bm{t}$ /* Undo flipped bits in output */
             end
            return Decoding failure
       end if
Function BP_with_oscillation($\bm{s},|\Phi|$):
[…]
       $\hat{\bm{e}}_{\text{prev}}\leftarrow\bm{0}$
      for $i=1$ to $i_{\max}$ do
[…]
             $\hat{\bm{e}}_{\text{prev}}\leftarrow\hat{\bm{e}}$
             if $\hat{\bm{e}}H^{T}=\bm{s}$ then
                   return True, $\hat{\bm{e}},\emptyset$
       end for
      /* Select top $|\Phi|$ most frequently flipped bits */
       $\Phi\leftarrow\textnormal{{top}}(\text{flip\_count},|\Phi|)$
       return False, $\hat{\bm{e}},\Phi$
Algorithm 1 BP decoding with Chase-like post-processing
Note that this concept is also widely adopted in classical decoding frameworks (Chase, 1972; Halford and Chugg, 2006; Dimnik and Be’ery, 2009; Wang et al., 2023). In such approaches, multiple modified inputs are generated and decoded independently, and the best result is selected according to the maximum likelihood criterion, typically based on minimum weight or log-likelihood score. Since each decoding attempt is independent, they can be executed in parallel, introducing minimal latency overhead. However, in the quantum setting, the structure of qLDPC codes introduces unique advantages. Due to the degeneracy and high distance of qLDPC codes, and the tendency of BP decoders to favor minimum-weight solutions, successful BP convergence rarely results in a logical error. This is partly because the classical codes $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$, defined by the parity-check matrices $H_{X}$ and $H_{Z}$, have low minimum distances (recall that the minimum distance of $\mathcal{C}_{X}$ is upper bounded by the row weight of $H_{Z}$, and vice versa). Even if BP converges to a non-optimal codeword within $\mathcal{C}_{X}$ or $\mathcal{C}_{Z}$, the resulting error is likely to be close to the minimal solution in terms of Hamming distance.
When considering the full quantum code defined by both $H_{X}$ and $H_{Z}$, the code distance is significantly higher than distances of $\mathcal{C}_{X}$ and $\mathcal{C}_{Z}$. This motivates our use of a speculative decoding strategy. In our approach, we omit the maximum likelihood selection step: Because of code degeneracy, any solution that satisfies the syndrome is likely to belong to the correct coset, particularly in the low-error regime and for high-distance codes. As a result, we simply return the first valid codeword that satisfies the syndrome among the parallel decoding attempts. This strategy reduces latency while preserving decoding performance.
5. Simulation Results
In this section, we evaluate the performance of the proposed decoder and compare it against the BP-OSD decoder. All BP decoders use the min-sum algorithm with an adaptive damping factor $\alpha=1-2^{i}$, where $i$ is the current number of iterations. The OSD method is OSD-CS in (Roffe et al., 2020), and for briefness, we use labels such as “BP1000-OSD10” to denote a decoder using BP with a maximum of 1,000 iterations followed by OSD-CS of order 10. Each data point is obtained by collecting at least 100 logical errors unless specified otherwise. As a result, the statistical uncertainty is sufficiently small, and error bars are omitted for clarity.
5.1. Code Capacity Model
In the code capacity error model, we decode using all test vectors of weight up to 1, which proves sufficient to achieve satisfactory logical error rates. In many of the codes we tested, the BP decoder already performs well under the code-capacity model, leaving limited room for improvement. However, there exist some exceptions. One such case is the $\llbracket 154,6,16\rrbracket$ coprime BB code introduced in (Wang and Mueller, 2024), where the min-sum decoder performs poorly despite the code’s high distance.
Fig. 4 shows the logical error rates of different decoders on the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code-capacity noise model. For the proposed decoder, the candidate set size is set to $|\Phi|=8$, resulting in a maximum of $50\times(8+1)=450$ BP iterations per decoding attempt. Considering that BP decoders can be run in parallel after the initial run, the latency is equivalent to 100 BP iterations. As shown, our proposed decoder significantly outperforms both the baseline BP and BP-OSD decoders, achieving lower logical error rates with fewer iterations and without requiring costly OSD post-processing. Additionally, both BP and BP-OSD exhibit an error floor at low physical error rates. Upon examining the decoding failures, we find that many are due to low-weight (e.g., weight-3) errors that were caused by trapping sets.
Figure 4. Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the code capacity model.
Another example is the $\llbracket 288,12,18\rrbracket$ BB code from (Bravyi et al., 2024). As shown in Fig. 5, the proposed decoder performs similar to the BP-OSD decoder while using fewer than $50\times(20+1)=1050$ iterations per decoding attempt and the latency is still 100 iterations, as we analyzed above.
Figure 5. Error rates of the $\llbracket 288,12,18\rrbracket$ BB code under the code capacity model.
5.2. Circuit-Level Noise Model
For the circuit-level noise model, errors are injected with uniform probability for gates and measurements. We use Stim (Gidney, 2021) to generate the detector error model and the corresponding parity-check matrix, where each row represents a detector event and each column corresponds to an error mechanism. Following the convention in prior literature, we perform $d$ rounds of syndrome extraction and define the logical error rate per round as
(11) LER Per $\text{LER Per Round}=1-(1-\text{LER})^{\frac{1}{d}},$
where LER is the logical error rate after $d$ rounds.
In this model, the number of error mechanisms is typically much larger than the number of qubits, resulting in very large parity-check matrices. Consequently, flipping a single bit as in the code capacity model is often insufficient for the BP decoder to converge. On the other hand, exhaustively decoding all test vectors with weight up to a threshold is computationally expensive. To address this, we adopt a sampling-based approach. Given a maximum test vector weight $w_{max}$, we randomly sample $n_{s}$ test vectors for each weight in $\{1,\dots,w_{max}\}$, resulting in a total of $n_{s}\times w_{max}$ test vectors per failed BP decoding attempt.
Figure 6 shows the logical error rates of different decoders on the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model. In this setting, the number of error mechanisms is approximately 8,000, which is significantly larger than the code length used in the code-capacity model. Therefore, we expand the candidate set size in the proposed decoder. Specifically, the proposed decoder uses up to 3,100 BP iterations (based on $w_{max}=6,n_{s}=5$) and achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder, which uses a maximum of 1,000 iterations and OSD order 10.
Figure 6. Error rates of the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model.
Figure 7. Error rates of the $\llbracket 154,6,16\rrbracket$ coprime-BB code under the circuit-level noise model.
Figure 7 shows the logical error rates of different decoders on the $\llbracket 144,12,12\rrbracket$ BB code under the circuit-level noise model. The proposed decoder uses up to 6,000 iterations (based on $w_{max}=6,n_{s}=10$. However, since all test vectors are decoded in parallel, the effective decoding latency corresponds to only 200 iterations. The proposed decoder achieves logical error rates that are slightly higher but still comparable to that of the BP-OSD decoder at lower physical error rate, which uses a maximum of 1,000 iterations and OSD order 10. In the higher physical error rate regime, the proposed decoder exhibits logical error rates that are higher than those of BP-OSD but still consistently lower than baseline BP decoding.
Figure 8. Error rates of the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model.
Figure 8 shows the logical error rates of different decoders on the $\llbracket 126,12,10\rrbracket$ coprime-BB code under the circuit-level noise model. The proposed decoder uses up to approximately 3,000 BP iterations to achieve a logical error rate comparable to that of the BP1000-OSD10 decoder. By increasing both $n_{s}$ and $w_{\max}$, we are able to further reduce the logical error rate to slightly below that of the BP-OSD decoder. However, this improvement comes at the cost of increased complexity, requiring up to 10,000 BP iterations.
5.3. Complexity and Parameter Selection
Next, we analyze the computational complexity of the proposed decoder. To ensure a fair comparison with baseline methods, we measure the average number of BP iterations under a serial execution model. Specifically, when the initial BP decoding fails, each test vector is decoded sequentially, and the total number of iterations is defined as the cumulative number of BP iterations required until the first successful decoding. This approach provides a conservative estimate of decoding cost, as it does not account for the inherent parallelism of our method, but allows for a meaningful comparison with standard decoders.
Figure 9 shows the growth in decoding complexity for the $\llbracket 144,12,12\rrbracket$ code as we target progressively lower logical error rates. For the BP decoder, we vary the maximum number of iterations to control decoding complexity. For the proposed decoder, we fix the maximum number of iterations per BP instance to 100 and vary $n_{s}$ (the number of test vectors sampled per weight), while keeping $w_{\text{max}}$ constant. This allows us to explore the trade-off between complexity and performance. The physical error rate is fixed at $3\times 10^{-3}$ under the circuit-level noise model. All data points are collected by simulating 10,000 shots for logical error rates above $10^{-3}$ and 100,000 shots for those below.
Across all decoders, we observe a linear region in which the number of BP iterations increases approximately linearly as the logical error rate decreases, up to a point where each curve drops off sharply, forming what we refer to as a cliff. This cliff marks a regime where the decoder can no longer reliably suppress logical errors within reasonable iteration limits. The proposed decoder consistently postpones this cliff compared to baseline BP, maintaining a lower iteration count at comparable logical error rates. Moreover, increasing $w_{max}$ increases the complexity but extends the linear region further and delays the start of the cliff, providing a tunable trade-off between decoding complexity and error suppression.
Figure 9. Complexity growth of different decoders. The parameter $|\Phi|$ is set to 50 for all proposed decoders. The number of iterations is calculated assuming serial execution.
6. Conclusion and Future Work
We introduced a fully parallelizable decoder based on belief propagation (BP). By leveraging speculative decoding and bit-flipping strategies guided by BP oscillation statistics, the proposed method achieves logical error rates comparable to BP-OSD, while significantly reducing computational complexity and avoiding costly Gaussian elimination. Extensive simulations show that our decoder performs exceptionally well under the code-capacity noise model across a range of bivariate bicycle codes. Under the more realistic circuit-level noise model, the decoder still delivers reasonable performance, though it requires a larger number of decoding trials to achieve comparable accuracy.
In future work, we aim to better understand the challenges posed by circuit-level noise and explore targeted improvements, such as more effective candidate selection, improved test vector sampling strategies, efficient decoder implementation, and enhancements to the inner BP decoder, in order to further improve decoding performance in practical fault-tolerant quantum computing systems.
Acknowledgment
The authors would like to thank John Stack for providing source code and for discussions on circuit-level noise simulation, and Timo Hillmann for valuable discussions regarding the BP-OSD decoder. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704, (Basic Energy Sciences, PNNL FWP 76274). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830. This work was also supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, MPS-2120757, CISE-2217020, and CISE-2316201 as well as DOE DE-SC0025384.
Appendix A Code Construction
In this section, we provide the construction parameters for the qLDPC codes referenced throughout the paper.
A.1. GB codes
Let $S_{l}$ be the shift matrix of size $l$, defined as
(12) $S_{l}=I_{l}>>1,$
where “>>much-greater-than$>$ >” denotes the right cyclic shift for each row in the matrix. For example,
(13) $S_{3}=I_{3}>>1=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}.$
Let $x=S_{l}$ here, The GB code can be defined by two polynomials, $a(x)$ and $b(x)$. These two polynomials can be represented by two matrices $A,B$ naturally, and the parity check matrices for the GB code are defined as
(14) $\begin{split}H_{X}&=[A|B]\\ H_{Z}&=[B^{T}|A^{T}].\end{split}$
The $\llbracket 254,28\rrbracket$ GB code used in this paper can be constructed by: $l=127$, $a(x)=1+x^{15}+x^{20}+x^{28}+x^{66}$, $b(x)=1+x^{58}+x^{59}+x^{100}+x^{121}$ as proposed in (Panteleev and Kalachev, 2021).
A.2. BB Codes
BB codes are constructed similarly to GB codes but with two variables. Let $x=S_{l}\otimes I_{m}$ and $y=I_{l}\otimes S_{m}$, where $\otimes$ denote Kronecker product, the BB codes can also be defined by two polynomials, $A=a(x,y)$ and $B=b(x,y)$. The parity check matrices for BB code are defined similarly as in GB codes.
The BB codes we used in the paper were proposed in (Bravyi et al., 2024). And the polynomials are shown in Table 1.
Table 1. BB Codes Used in Simulations
$l$ $m$ $a(x,y)$ $b(x,y)$ $\llbracket n,k,d\rrbracket$
6 6 $x^{3}+y+y^{2}$ $y^{3}+x+x^{2}$ $\llbracket 72,12,6\rrbracket$
12 6 $x^{3}+y+y^{2}$ $y^{3}+x+x^{2}$ $\llbracket 144,12,12\rrbracket$
12 12 $x^{3}+y^{2}+y^{7}$ $y^{3}+x+x^{2}$ $\llbracket 288,12,18\rrbracket$
A.3. Coprime-BB Codes
The coprime-BB code we used were found in (Wang and Mueller, 2024). Let $\pi=xy$, where $x$ and $y$ are defined the same as in BB codes. The coprime-BB code we tested can be constructed as shown in Table 2.
Table 2. Coprime-BB Codes Used in Simulations
$l$ $m$ $a(\pi)$ $b(\pi)$ $\llbracket n,k,d\rrbracket$
7 9 $1+\pi+\pi^{58}$ $1+\pi^{13}+\pi^{41}$ $\llbracket 126,12,10\rrbracket$
7 11 $1+\pi+\pi^{31}$ $1+\pi^{19}+\pi^{53}$ $\llbracket 154,6,16\rrbracket$
Appendix B ”Good” Codes for BP
This section presents several codes that demonstrate good performance under BP decoding, along with their corresponding logical error rates. Since both BP-OSD and the proposed decoder act as post-processing techniques, they are only invoked when the BP decoder fails to converge. As a result, for the codes shown below, where BP alone achieves high success rates, we observe similar overall performance across all decoding strategies.
B.1. Code Capacity Model
Figure. 10(a-d) shows the performance of various decoders on several codes under the code-capacity noise model. In these cases, the baseline BP decoder already achieves logical error rates comparable to those of the BP-OSD decoder. In these cases, post-processing yields only marginal improvements.
Figure 10. The logical error rates of different codes under code capacity error model. For (a): $\llbracket 72,12,6\rrbracket$ BB code, BP and BP-OSD have the same error when we set the seed to the same.
B.2. Circuit Level Noise Model
For the $\llbracket 72,12,6\rrbracket$, the BP and BP-OSD decoder have similar performance, as shown in Fig. 11.
Figure 11. Error rate comparison on the $\llbracket 72,12,6\rrbracket$ under circuit-level noise model.